Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Tuesday, December 15, 1998

The following is a brief summary of the environmental monitoring activities that were completed on December 15, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

12/15/98 (Tuesday)

Segment 1 report for 12/15/98: Calibration continued at Emidio Pump Station. A sinkhole at ~EMP 3.76 on Aqueduct Access Rd. A was being. At Stream Crossing #4 (EMP 9.03) tie-in of the pipe in the bore to the main line occurred on the east side from EMP 9.04 to 9.07. Backfill was in progress at the Tecuya Creek/I-5 bore (EMP 9.67). !5 feet of casing has been installed at the Stream Crossing #9 bore (EMP 15.35), progress here is slow due to an abundance of rock. On Grapevine Rd., conduit has been placed and backfill was in progress from EMP's 16.27 to 16.07. Pullboxes are being installed at 15.92 and 16.14. Electrical work continued at Grapevine Pump Station, rock was being spread around the site. Bore equipment was installed in the Tejon Bench/I-5 bore (EMP 16.49). On Tejon Bench shading was in progress from EMP 16.65 to 16.57. A segment that was skipped from EMP 17.73 to 17.68 has been trench, pipe has been placed and welding is under way. Excavation has resumed at EMP 18.03, backfill was occurring to 18.06. On Lebec Rd., excavation occurred to EMP 21.64. Conduit was placed and backfill occurred to EMP 21.61. No work was done at the Cuddy Creek Tributary bore (EMP 25.47).

Segment 2B (Castaic to Sylmar): Report for 12-15-98. Fiber optic duct was tested then cable pulled at the Santa Clara river. In the PGA Ranch section between Valencia Blvd and McBean Pkwy crews repaired a section of the right of way damaged during a rain event. Meanwhile electrical crews progressed with the wire harness at Pico Cyn Rd and crews backfilled the repair pit south of Lyons Ave. General clean up and removal of debris, shoring and materials continued at Weldon Cyn and paving crews completed work at the vault site near EMP 72.4 In addition, paving crews capped the roadway between Sierra Hwy and Sunshine Cyn. Segment 2C(Templin Hwy to Castaic): Crews capped the I-5 road bore with concrete to allow filling of the pipe with clay as requested by Caltrans.

Segment 3A & 3C (Sylmar to Los Angeles): Report for 15 December 1998: 3A: Backfill crews moved forward with closing up the ditch at San Fernando and Hollywood Way Bridges. Welding, x-ray, and coating focused on the north side tie in to the Tuxford bore. Crews also worked in this area screening backfill material. 3C: Fiber optic crews worked between Grapevine Creek and Polk Street today pulling fiber. They were joined by a sub contractor for the City of Los Angeles, who worked on pulling their 132 fibers. The crew at Hubbard Avenue neared completion of shading the pipe, as well as placing a pull box on the south side of Hubbard for the FOC. Some slurry was also poured in this location. Worked on both the aqueduct crossing and high speed channel cover progressed today. Welding continued to set up the 6" pipe for the Hydrotest filling and discharge upcoming.

Inspection Summar

In the past 24 hours, 3 inspection event(s) were reported for 3 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page